OpenStudy (anonymous):

x^2+y^2=26 y-x=-6 Solving system of non linear equations

OpenStudy (anonymous):

y=x+6 x^2+(x+6)^2 x^2+x^2+36=26 2x^2+36=6 2x^2=-10 x^2=5

OpenStudy (anonymous):

^That's right. Just plug in X-6 for your Y and solve.